Previous | Calendar |
October |
2003 |
Next | |||
---|---|---|---|---|---|---|---|
Sunday | Monday | Tuesday | Wednesday | Thursday | Friday | Saturday | |
28JD271 |
29JD272 |
30JD273Site Visit (SpB, BB): Details , Photos (17) |
1JD274 |
2JD275Site Visit (SpB): Details , Photos (15) West Robot Flasks Swapped Removed MUXII from TC cr23x logger (didn't change csi program) |
3JD276Site Visit (SpB,BB): Details ~11AM, TGA Dewar Filled (SpB) |
4JD277 |
|
5JD278 |
6JD279 |
7JD280Site Visit (SpB,AT,RM,AS, JS,DL,TD,KK,NP): Details ~9:30AM, TGA Dewar Filled (SpB) Replaced East Robot (Andy S.) Started moving soil sensor (disconnected soil logger) (AT) |
8JD281 |
9JD282Site Visit (SpB,AT,RM,CXY): Details ~11AM, TGA Dewar Filled (SpB) West Robot Flasks Swapped Moved soil sensors (heat flux plates, temp, moisture sensors), restarted soil logger (using same program) |
10JD283 |
11JD284 |
|
12JD285 |
13JD286Site Visit (SpB,BB): Details ~10AM, TGA Dewar Filled (BB) Ameriflux Meeting |
14JD287Ameriflux Meeting |
15JD288Ameriflux Meeting |
16JD289Site Visit (AT,DB): Details Ameriflux Meeting Laser Teardown |
17JD290Site Visit (SpB,AT,RM,DB,BB, CXY,LSD,DL,JH,S): Details Laser Teardown Hydra Installed East/West Robots both had done==2 (not changed), checked by Dave Bowling |
18JD291 |
|
19JD292 |
20JD293 |
21JD294Site Visit (SpB): Details At 13:15 MST refurbished LiCor 6262 up and running. |
22JD295Site Visit (SpB, AT, RM): Details --data system shutdown at ~4am--- (restarted at about 11am) |
23JD296Site Visit (SpB): Details Robot Loggers Reset. |
24JD297 |
25JD298Site Visit (SpB): Details |
|
26JD299--data system shutdown at ~4am--- (restarted at about 9am) |
27JD300 |
28JD301--data system shutdown at 23:38 MST--- (restarted on 9/29 @ 8:30am) |
29JD302Site Visit (SpB, AT): Details , Photos (7) Robot Loggers Reset. Note: Overland Fire Burning in LeftHand Canyon. Additional TCs added to soil logger..new program started: soilsnow.csi TC logger program changed to: Tc23x03.csi (removed muxII/III, added IRT (IRT not working)) |
30JD303Brief data system shutdown at 13:00 MST--- (due to a short power outage, restarted at 13:10 MST) |
31JD304 |
1JD305Site Visit (SpB): Details , Photos (55) |